OpenStudy (anonymous):

solve the equation or inequality. 2m+7=17 please show work

OpenStudy (anonymous):

2m+7=17 2m=10 for this step you subtracted 7 from the left side and right sides m=5 for this step you divided both sides by two to get your final answer

OpenStudy (anonymous):

2m+7=17 -7=-7 2m=10 m=5 you have to subtract 7 from both sides, your left with 2m=10 divide 2 from both sides and you will get m=5